2x^3-10x^2=0
show work

Answer by Fermat(127)

2x^3-10x^2=0
Take out a common factor of x^2 first of all.
x^2(2x - 10) = 0
Therefore, x^2 = 0, 2x-10 = 0
or,
x = 0, x = 5
============